OpenStudy (anonymous):

What exponential function is the best fit for the data in the table? x f(x) 1 -4 3 -1 4 3 f(x) = 4(3)x - 1 + 4 f(x) = 4(3)x - 1 - 4 f(x) = one fourth(3)x - 1 + 4 f(x) = one fourth(3)x - 1 - 4

OpenStudy (anonymous):

i plugged in the x's but it just doesn't make sense

OpenStudy (anonymous):

f(x) = 4(3)^(x - 1) + 4 f(x) = 4(3)^(x - 1) - 4 f(x) = 1/4(3)^(x - 1) + 4 f(x) = 1/4(3)^(x - 1) - 4

OpenStudy (anonymous):

@wio @precal

OpenStudy (anonymous):

What subject is this?

OpenStudy (anonymous):

algebra 11

OpenStudy (anonymous):

2

OpenStudy (anonymous):

First of all, the function is increasing, so the base of the exponent should be \(>1\) ok?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

All exponential functions can be of the form :\[ f(x) = b^x+c \]

OpenStudy (anonymous):

In our case, we say \(b>1\).

OpenStudy (anonymous):

but we already have our b

OpenStudy (anonymous):

dont we?

OpenStudy (anonymous):

whoops \[ f(x) = ab^x+c \]

OpenStudy (anonymous):

What is \(b\)?

OpenStudy (anonymous):

i believe from the answer choices it would be 3

OpenStudy (anonymous):

When \(x = 0\), then we have \[ ab^0+c = a+c \]

OpenStudy (anonymous):

true

OpenStudy (anonymous):

It looks like they have \[ ab^{x-1}+c \]So in this case, \(x=1\) would be where our \(x=0\) would have been.

OpenStudy (anonymous):

pretty much but it still does not match the chart

OpenStudy (anonymous):

What is closest?

OpenStudy (anonymous):

ooooh am i suppose look for the closest?

OpenStudy (anonymous):

well the closest is D

OpenStudy (anonymous):

D gets -3.75

OpenStudy (anonymous):

hmmmm........

OpenStudy (anonymous):

nothing else actually works A = 8 B = 0 C = 4.25 D = -3.75

OpenStudy (anonymous):

@wio are you there

OpenStudy (anonymous):

well thanks for your help wio you helped me get a 100%

OpenStudy (anonymous):

legit

OpenStudy (anonymous):

gracias